OpenStudy (anonymous):

If x^2=25-y what is the value of d^(2)y/dx^(2) at the point (3,4)

OpenStudy (gorv):

\[x^2=25-y\] \[y=25-x^2\] \[dy/dx=0-2x\] \[d^2y/dx^2 = -2\] so at (3,4 ) d^(2)y/dx^(2)=-2